OpenStudy (anonymous):

It's asking which equation you would use to isolate a variable. Explain your reasoning. Problem: x=8y-3 3x-4y=1

OpenStudy (anonymous):

The equation already has X solved for you, so plugging X from eq.1 to eq2 yields 3(8y-3)-4y =1. Distribute. 24y-9-4y=1. Isolate Y. 20y=10, Y = 10/20 = .5 You can then plugin Y in the other equation to solve for X. x = 8(.5)-3 x = 4-3 x = 1

OpenStudy (anonymous):

okay. And i need help on another problem. Its asking to solve the linear system by using substitution. -x+2y=-6 8x+y=31

OpenStudy (anonymous):

Like the other problem, you want to isolate a variable. x is easier to isolate again in eq1. x - 2y + 6 substitution into eq2. 8(2y+6) + y = 31 16y + 48 + y = 31 17y = -17 y = -1 Plugin Y into eq. 1 to get X. X - (2*-1) + 6 = -4 y = -1 x = 4